The Department of Agriculture of Economica (DAE) administers the floor price of milk at $4 per pound of milk. To support the price of milk at the price floor, the DAE had to buy up the surplus.

Q = 120 – 20P (Market demand)

Q = 20P (Market supply)

ii) Explain floor price and the purpose for the government to implement the policy.

iii) Use the equations to determine the quantity demanded and quantity supplied of milk at

the floor price. Explain the market outcomes of the floor price.

iv) Calculate the consumer surplus and producer surplus in the absence of a price floor.

v) Calculate the consumer surplus and producer surplus with the price floor at $4 per pound of milk.

vi) Use the results in parts (iv) and (v) to explain the impact of floor on the economic welfare of consumers and producers.

vii) Calculate the deadweight loss of floor price policy and explain the result.

viii) How much money does the DAE spend on buying up surplus milk?


1
Expert's answer
2021-12-29T12:30:23-0500

Solution:

ii.). Floor price refers to price control or limit imposed by the government or a group on how low a price can be charged for a product, good, commodity, or service. To be effective, a price floor must be higher than the equilibrium price.

 

iii.). At equilibrium: Qd = Qs

120 – 20P = 20P

120 = 20P + 20P = 40P

120 = 40P

P = 3

The market equilibrium price = $3

Substitute to derive the market quantity:

Q = 120 – 20P = 120 – 20(3) = 120 – 60 = 60

The market equilibrium quantity = 60

The floor price of $4 is above the equilibrium price of $3.

When a price floor is set above the equilibrium price, the quantity supplied exceeds the quantity demanded, resulting in excess supply or surpluses.

 

iv.). Consumer surplus = ½ "\times" 60 "\times" (6 – 3) = 90

Producer surplus = ½ "\times" 60 "\times" (3 – 0) = 90

 

v.). Consumer surplus = ½ "\times" 40 "\times" (6 – 4) = 40

Producer surplus = ½ "\times" 40 "\times" (3 – 2) = 20

 

vi.). The floor price will decrease the economic welfare of both consumers and producers.

 

vii.). Deadweight loss = ½ "\times" (4 – 2) "\times" (60 – 40) = 20

This is the loss incurred due to the floor price introduced in the market.

 

viii.). DAE spend on buying up surplus milk = 20 "\times" 3 = $60